# Localization landscape of optical waves in multifractal photonic membranes

### 26 Jan 2024 v1

@@ 5.2 Multifractal photonic membrane fabrication
A sketch of the process flow is shown in Figure 7(a). First a negative tone resist is spun atop of the grown material to achieve a thickness of $\approx$100 nm followed by soft bake. A thin, conductive coating is then spun atop of the resist to dissipate charging during the EBL process. Nano-hole patterns are then written onto the photoresist via EBL using a 30kV source (Zeiss Supra40VP). To develop the patterns, the sample is rinsed with DI water to remove the protective conductive coating and then submerged in a 4:1 ratio of developer and DI water for 55-60 seconds and blow dried with N2 gas. A brief plasma ashing step in a pure oxygen environment is performed in order to remove residual resist left over from the development stage. This step is crucial in achieving reliable liftoff. A 20 nm chromium mask is then evaporated onto the sample (Step 3) and submerged in acetone for lift-off. A dry etch of the exposed nano-holes is performed using anisotropic reactive ion etching and finally a wet chromium etch is performed to reveal the final nano-hole structure. Representative scanning electron microscope (SEM) images of a multifractal sample are shown in Figure 7(b) along with dark-field optical images of all multifractal structures fabricated are shown in Figures 7(c-f). All patterns were written onto a square 100 $\mu m^{2}$ area with each hole having the desired radius of 100nm as confirmed by SEM imaging. In order to facilitate the comparison of the measured and simulated optical modes across the investigated spectral range, both monofractal and multifractal membranes were fabricated with $\overline{d}_{1}/\lambda=0.62$ at the operation wavelength $\lambda=700\,\mathrm{nm}$.
@@ 5.3 Experimental characterization of multifractal resonances
The out of plane scattered radiation is collected by a high NA objective (50x) and an image of the modes is focused onto a camera (Thorlabs CS126MU) using a tube lens. We first measured the leaky modes of the fabricated monofractal structure as a function of wavelength from 450nm-750nm. In Figure 8 we show representative images of the spatial distributions of the optical modes of the monofractal structure excited using TE-polarized light at different wavelengths. We observe clearly distinct spatial distributions of the modes when excited at different wavelengths. Moreover, all mode patterns from these data demonstrate spatial localization within small particles’ clusters with highly fluctuating intensity profiles that are qualitatively similar to the numerical predictions from the landscape theory. Furthermore, we also observed that the mode patterns excited at the smallest wavelengths exhibit a stronger degree of spatial localization and are mostly confined inside the geometrical support of the monofractal structure, as also evident from the numerical results previously shown in Figure 4. This characteristic modal clustering behavior around smaller regions of space of the monofractal geometry is also manifested by the calculated modes in Figure 4 and is in fact expected for self-similar structures in which scattering resonances are driven by coupling effects within locally-symmetric particles’ clusters occurring at multiple scales [34, 41].
@@
As a comparison, we also investigated the optical resonances of the multifractal structure generated with the probability vector $p=[1,0.75,0.5,0.25]$ and a spectral parameter identified using the landscape theory, enabling a direct experimental comparison of the leaky-mode structures of fractals and multifractal systems at minimal computational cost. As shown in in Figure 9, the observed mode patterns of the multifractal structure change according to the incident wavelengths and display highly fluctuating spatial profiles. However, in contrast to the case of the monofractal structure, the multifractal one demonstrates spatially extended modes distributed across the entire geometrical support, in good qualitative agreement with the simulation results shown in Figure 5. Moreover, due to the more spatially inhomogeneous nature of this multifractal structure [20], the measured modes spread across the structure with no simple trend in the degree of spatial localization. These results, demonstrate the distinctive localization behavior of the scattering resonances supported by large-scale fractal and multifractal photonic membranes designed by the efficient localization landscape approach. Future work will focus on the quantitative analysis of the spectral measures, such as the integrated density of states [17], of the investigated multifractal optical systems.
@@ 6 Conclusion
Acknowledgments L.D.N. acknowledges support from the National Science Foundation (ECCS-2015700, ECCS-2110204) and insightful discussions with Marcel Filoche.

### 05 Mar 2024 v2

@@ 5.2 Multifractal photonic membrane fabrication
Figure 8: (a) Fabrication process flow via EBL and anisotropic RIE. (b) Representative SEM image of the fabricated multifractal pattern. The inset is a close up view of the same sample. The scale bars in the SEM image and inset are 4 $\mu m$ and 600 nm respectively. (c-f) Dark field scattering images of monofractal and multifractal structures. Each array consists of $\approx$15,000 points written over a 100 $\mu m^{2}$ area. The initial probability vector used to generate each pattern is reported above each image.
@@ 5.3 Experimental characterization of multifractal resonances
The out of plane scattered radiation is collected by a high NA objective (50x) and an image of the modes is focused onto a camera (Thorlabs CS126MU) using a tube lens. We first measured the leaky modes of the fabricated monofractal structure as a function of wavelength from 450 nm-750 nm. In Figure 9 we show representative images of the spatial distributions of the optical modes of the monofractal structure excited using TE-polarized light at different wavelengths. We observe clearly distinct spatial distributions of the modes when excited at different wavelengths. Moreover, all mode patterns from these data feature spatial localization within small clusters of particles with highly fluctuating intensity profiles that are qualitatively similar to the numerical predictions from the landscape theory. This characteristic modal clustering behavior around smaller regions of space of the monofractal geometry is also manifested by the calculated modes in Figure 5 and is expected for self-similar structures in which scattering resonances are driven by coupling effects within locally-symmetric clusters distributed at multiple length scales [37, 44].
As a comparison, we also investigated the optical resonances of the multifractal structure generated with the probability vector $p=[1,0.75,0.5,0.25]$ and a spectral parameter identified using the landscape theory, enabling a direct experimental comparison of the leaky-mode structures of fractals and multifractal systems at minimal computational cost.
@@
As shown in in Figure 10, the observed mode patterns of the multifractal structure display highly fluctuating spatial profiles distributed across the entire geometrical support, in qualitative agreement with the simulation results shown in Figure 6. Moreover, due to the spatially non-homogeneous nature of the investigated multifractal structure, the measured modes display no evident trend in the degree of their spatial localization when varying the wavelength, which is indicative of a broader spectrum of localized resonances compared to monofractals [20]. These results, demonstrate the distinctive localization behavior of the scattering resonances supported by large-scale fractal and multifractal photonic membranes designed by the efficient localization landscape approach. Future work will focus on the quantitative analysis of the spectral measures of the investigated multifractal optical systems, such as their integrated density of states [17].
@@ 6 Conclusion
Acknowledgments L.D.N. acknowledges support from the National Science Foundation (ECCS-2015700, ECCS-2110204) and insightful discussions with Marcel Filoche and Sergey Skipetrov.